# Enhancing Knowledge Distillation of Large Language Models through Efficient Multi-Modal Distribution Alignment

### v1

@@ Abstract
Knowledge distillation (KD) is an effective model compression method that can transfer the internal capabilities of large language models (LLMs) to smaller ones. However, the multi-modal probability distribution predicted by teacher LLMs causes difficulties for student models to learn. In this paper, we first demonstrate the importance of multi-modal distribution alignment with experiments and then highlight the inefficiency of existing KD approaches in learning multi-modal distributions. To address this problem, we propose Ranking Loss based Knowledge Distillation (RLKD), which encourages the consistency of the ranking of peak predictions between the teacher and student models. By incorporating word-level ranking loss, we ensure excellent compatibility with existing distillation objectives while fully leveraging the fine-grained information between different categories in peaks of two predicted distribution. Experimental results demonstrate that our method enables the student model to better learn the multi-modal distributions of the teacher model, leading to a significant performance improvement in various downstream tasks [^1] Our code is available at https://github.com/Pty72/RLKD..
@@
Tianyu Peng  and Jiajun Zhangthanks: Corresponding Author. School of Artificial Intelligence, University of Chinese Academy of Sciences Institute of Automation, Chinese Academy of Sciences Wuhan AI Research pengtianyu2022@ia.ac.cn, jjzhang@nlpr.ia.ac.cn
@@ 1 Introduction
In recent years, large language models (LLMs, Brown et al. 2020; Zeng et al. 2023; OpenAI 2023; Touvron et al. 2023; Yang et al. 2023; Jiang et al. 2024) have demonstrated their great power to solve natural language processing tasks. Researches Kaplan et al. (2020); Wei et al. (2022) show that language models tend to perform better as the number of parameters increases. However, the training and deployment of large scale models involve high costs, coupled with less usability and flexibility. Therefore, model compression techniques Zhu et al. (2023) for LLMs draw more and more attention.
Knowledge distillation (KD, Hinton et al. 2015) is one of the representative approaches for model compression. It achieves efficient knowledge transfer to smaller student models by conveying the complete probability distribution output of teacher models. Therefore, designing better distillation objectives that make it more efficient for student models to learn the overall probability distribution of teacher models is the focus in KD research.
Particularly for LLMs, the lengthy and complex probability distribution causes more learning difficulties. Due to the diversity of natural language, the predicted probability distribution of LLMs are often multi-modal (quantitative analysis results are shown in Appendix A), which contains multiple potential correct predictions for a given input. Consequently, improving the learning of multi-modal distribution becomes the focus of LLMs’ KD.
To achieve this goal, previous studies have already explored several distillation objectives for LLMs. Conventional KD Hinton et al. (2015) uses forward Kullback-Leibler divergence (KL) as the optimization objective. However, KL predisposes to the mode-averaging problem Wen et al. (2023), whereby student models tend to learn too smooth distributions (as shown in Figure 1). Therefore, more recent work Gu et al. (2023); Tan et al. (2023) has employed reverse KL (RKL) instead of KL, claiming that this optimization objective can better focus on peak predictions. Nevertheless, optimizing RKL tends to get an overconcentration of the probability predictions of the student model in some specific intervals (as illustrated in Figure 1). Then, Wen et al. (2023) propose using symmetric divergences as the distillation objective to alleviate mode problems caused by KL and RKL. In contrary, Wu et al. (2024) verify through theory and experiments on toy data that KL and RKL do not suffer from the above issues, and they instead share the same optimization objective in KD of LLMs.
@@
In order to address the above issues, we propose in this paper Ranking Loss based Knowledge Distillation (RLKD) for LLMs. We first verify the relationship between multi-modal predictions and model performance, and experimentally demonstrate the problems of existing distillation objectives in fitting multi-modal distributions. In response to the identified problems, we introduce the word-level ranking loss, which is based on Spearman’s rank correlation coefficient (SRCC), to optimize the degree of consistency in the order of peak predictions between the teacher and student models. In particular, we transform the learning of multi-modal distributions during the KD process into the learning of the top-k sampling Holtzman et al. (2020) order. Through ranking loss, we ensure excellent compatibility with existing distillation objectives while fully leveraging the fine-grained information between different categories predicted by two distribution peaks. Additionally, we verify and demonstrate through real-scenario experiments the impact of introducing ranking loss into KD on the learning ability of student models for multi-modal distributions.
Experimental results indicate that the quality of multi-modal predictions is closely related to the performance of the model, while existing distillation objectives lack the ability to fit multi-modal distributions effectively. Consequently, our proposed method effectively enhances the student model’s learning ability to predict multi-modal distributions during the distillation process and exhibits good compatibility with existing distillation objectives. We also validate the ranking loss on diverse datasets from multiple tasks, showing significant improvements in downstream KD tasks.
@@ 2.1 KD of LLMs
Nowadays, many LLMs are no longer open source due to commercial and other considerations. Therefore, based on the open-source nature of the model, KD of LLMs is frequently categorized into white-box KD Gu et al. (2023); Tan et al. (2023) for open-source LLMs Touvron et al. (2023); Yang et al. (2023); Jiang et al. (2024) and black-box KD Zhou et al. (2023); Chen et al. (2024) for closed-source LLMs Brown et al. (2020); OpenAI (2023).
@@ 2.2 White-box KD Objectives
Unlike the above, Wu et al. (2024) argue that the predicted distribution of LLMs does not meet the conditions of continuity and standard Gaussian distribution. They theoretically and practically demonstrate that KL and RKL actually share the same optimization objective in LLMs’ KD. Additionally, they point out that KL and RKL follow different optimization paths, with one fitting from the head part first and the other from the tail part first, and propose dynamically combining KL and RKL into Adaptive Kullback-Leiber (AKL).
Inspired by these researches, we aim at exploring a KD objective that yields better multi-modal learning ability. However, we first need to validate the following two questions:
@@ 3.1 Metrics
The purpose of enhancing the ability to learn multi-modal distributions is to make the peak predictions of student and teacher models closer. Therefore, we first need to define the criteria for evaluating the similarity of the peak predictions of two models. Since top-k sampling is the most common sampling strategy in language models, and peak predictions well correspond to the results of top-k sampling. Hence we transform the measurement of the consistency of multi-modal predictions into the measurement of the consistency of top-k sampling results.
@@ 3.3 Validation for Existing KD Objectives
Analysis Based on the results in Figure 3, existing distillation objectives show no significant disparity in their impact on the ability to learn multi-modal distributions. This further confirms Wu et al.’s (2024) view that KL and RKL share the same optimization objective in KD of LLMs. But more importantly, even after 20 epochs of training, student models still exhibit deficiencies in learning multi-modal distributions under the existing distillation objectives. Therefore, further exploration is necessary to identify distillation metrics that can enhance the model’s capability to learn multi-modal distributions.
@@ 4 Method
In this work, we enhance the learning of peak predictions in KD of LLMs by introducing a new optimization objective of word-level ranking loss. The new optimization objective focuses on the prediction order of high probabilities between student and teacher models, enabling student model to match the teacher model on critical predicted categories.
@@
Our approach allows that during the optimization process, the calculation of peak predictions is not limited to comparisons within a single category. Instead, as the red lines shown in Figure 4, $Q(i)$ needs to be compared with $Q(j)$ and $Q(k)$ based on the ranking position of $P(i)$ in the teacher’s predictions to minimize the ranking loss.
@@ 5.1 Baselines
Adaptive Kullback-Leiber divergence (AKL, Wu et al. 2024) For the calculation process of AKL, we use the same experimental setup as in the original paper. We set the hyperparameter $\mu$ as 0.5 and the gap function $\epsilon(p(z),q(z))=|p(z)-q(z)|$.
@@ 5.2 Datasets and models
databricks-dolly-15k Conover et al. (2023) A directive fine-tuning dataset covering various tasks. We randomly select 14,000 samples for the training set and 800 samples for the test set. We use ROUGE scores Lin (2004) as the evaluation metric, abbreviated as "R-1", "R-2" and "R-L".
@@
For all KD tasks, we employ Llama-2-7B Touvron et al. (2023) as the teacher model and Tinyllama-1.1B Zhang et al. (2024) as the student model. Prior to distillation, we fine-tuned the teacher model on the respective datasets to adapt it to the tasks. Except in the GSM8K task, we directly use gsm8k-rft-llama7b2-u13b model released by Yuan et al. (2023) due to its excellent performance.
@@ 5.3 Results in the Pre-training Task
Training With the same experimental setup as in Section 3.3, we test the effectiveness of these five distillation objectives combined with ranking loss. For ranking loss, we align predictions of the top 15 between student and teacher models. We employ a fixed ratio for loss allocation, as $\mathcal{L}_{\text{total}}=2\cdot\mathcal{L}_{\text{Ranking}}+\mathcal{L}_{\text{logits}}$. The choice of ranking ranges and the allocation of losses is discussed in Appendix D. The experimental results are presented in Table 1, while more detailed training information can be found in Appendix B. We also present extra experiments in Appendix E to verify the generalization ability of our method.
Results We present visualized results in Figure 5, which more intuitively show the improvement in peak predictions consistency of ranking loss compared to the original distillation objectives. The introduction of ranking loss significantly improves the efficiency of aligning multi-modal predictions between the student and teacher models during the distillation process. Furthermore, for the five different distillation objectives, ranking loss demonstrates stable and effective improvements, showcasing its compatibility with existing objectives.
5.4 Results in Downstream Tasks
@@
Table 2: Experimental results on test set of downstream tasks. "+R" represents that we have added an additional fixed-ratio ranking loss.
We verify the effect of ranking loss on different downstream tasks.
@@
Results The experimental results demonstrate a significant improvement in the performance of downstream tasks when introducing the ranking loss. Our method improves the performance of different KD objectives on various tasks. Especially, when only using the ranking loss, our method learns only the peak predictions, which account for 0.01% of the total categories, yet surpasses most existing distillation objectives on multiple tasks. This not only demonstrates the importance of aligning peak predictions, but also showcases the outstanding performance of our method in peak predictions alignment.
@@ 6.1 Ablation Study
Based on the respective results in Table 1 and Table 2, we can conclude that the improvement in the ability to learn multi-modal distributions and the enhancement in downstream task performance are indeed attributed to the introduction of ranking loss.
Furthermore, we can observe in Table 1 that although the improvement of the top prediction accuracy is modest by introducing the ranking loss, the peak prediction alignment at other positions is significantly improved.
However, in downstream tasks, we use a greedy decoding strategy, which should not exhibit better performance when there is no significant improvement in the consistency of the top prediction. In fact, the reason is that in downstream tasks, we make the student model learn task-related prediction patterns more efficiently by enhancing its ability to learn from multi-modal distributions. Therefore, our method can also achieve better performance with the same number of training steps.
Based on the above analysis, we can more confidently get the conclusion that ranking loss primarily improves the fitting of multi-modal distributions, with modest impact on the alignment of top-one predictions. Therefore, the existing objectives without the addition of ranking loss can be regarded as an ablation of the ability to learn multi-modal distributions, leading to worse results. This further shows the importance of aligning multi-modal predictive distributions for KD in LLMs.
@@ 6.2 Case study
We conduct a case study within this interesting phenomenon in Appendix G to analyze to further demonstrate the importance of peak predictions learning.
@@ 7 Conclusion
In this paper, we propose ranking loss based knowledge distillation, a new objective function that improves the efficiency of aligning peak predictions of student and teacher models during white-box KD. We verify the importance of aligning multi-modal distributions through experiments and highlight the inefficiency of existing KD objectives in learning multi-modal distributions. Most importantly, we propose a word-level ranking loss to the existing KD objectives for more efficient alignment of multi-modal distributions. Extensive experiments demonstrate that our method effectively improves the multi-modal distribution alignment between teacher and student models, leading to significant performance gains in different downstream tasks.

### v2

@@ Abstract
Knowledge distillation (KD) is an effective model compression method that can transfer the internal capabilities of large language models (LLMs) to smaller ones. However, the multi-modal probability distribution predicted by teacher LLMs causes difficulties for student models to learn. In this paper, we first demonstrate the importance of multi-modal distribution alignment with experiments and then highlight the inefficiency of existing KD approaches in learning multi-modal distributions. To address this problem, we propose Ranking Loss based Knowledge Distillation (RLKD), which encourages the consistency of the ranking of peak predictions between the teacher and student models. By incorporating word-level ranking loss, we ensure excellent compatibility with existing distillation objectives while fully leveraging the fine-grained information between different categories in peaks of two predicted distribution. Experimental results demonstrate that our method enables the student model to better learn the multi-modal distributions of the teacher model, leading to a significant performance improvement in various downstream tasks. [^1] Our code is available at https://github.com/Pty72/RLKD.
@@
Tianyu Peng1,2,3  and Jiajun Zhang1,2,3,4thanks: Corresponding Author. 1School of Artificial Intelligence, University of Chinese Academy of Sciences 2Institute of Automation, Chinese Academy of Sciences 3Wuhan AI Research, 4Shanghai Artificial Intelligence Laboratory, Shanghai, China pengtianyu2022@ia.ac.cn, jjzhang@nlpr.ia.ac.cn
@@ 1 Introduction
In recent years, large language models (LLMs, Brown et al. 2020; Zeng et al. 2023; OpenAI 2023; Touvron et al. 2023; Yang et al. 2023; Jiang et al. 2024) have demonstrated their great power to solve natural language processing tasks. Existing research Kaplan et al. (2020); Wei et al. (2022) shows that language models tend to perform better as the number of parameters increases. However, the training and deployment of large scale models involve high costs, coupled with less usability and flexibility. Therefore, model compression techniques Zhu et al. (2023) for LLMs are drawing more and more attention.
Knowledge distillation (KD, Hinton et al. 2015) is one of the representative approaches for model compression. It facilitates efficient knowledge transfer to smaller student models by using the full probability distribution output from teacher models as a guiding signal during optimization. Therefore, designing better distillation objectives that make it more efficient for student models to learn the overall probability distribution of teacher models is the focus in KD research.
Particularly for LLMs, the lengthy and complex probability distribution causes more learning difficulties. Due to the diversity of natural language, the predicted probability distribution of LLMs is often multi-modal (quantitative analysis results are shown in Appendix A), which contains multiple potential correct predictions for a given input. Consequently, improving the learning of multi-modal distribution becomes the focus of LLMs’ KD.
To achieve this goal, previous studies have already explored several distillation objectives for LLMs. Conventional KD Hinton et al. (2015) uses forward Kullback-Leibler divergence (KL) as the optimization objective. However, KL predisposes to the mode-averaging problem Wen et al. (2023), whereby student models tend to learn too smooth distributions (as shown in Figure 1). Therefore, more recent work Gu et al. (2023); Tan et al. (2023) has employed reverse KL (RKL) instead of KL, claiming that this optimization objective can better focus on peak predictions. Nevertheless, optimizing RKL tends to get an overconcentration of the probability predictions of the student model in some specific intervals (as illustrated in Figure 1). Then, Wen et al. (2023) propose using symmetric divergences as the distillation objective to alleviate mode problems caused by KL and RKL. In contrary, Wu et al. (2024b) verify through theory and experiments on toy data that KL and RKL do not suffer from the above issues, and they instead share the same optimization objective in KD of LLMs.
@@
In order to address the above issues, we propose in this paper Ranking Loss based Knowledge Distillation (RLKD) for LLMs. We first verify the relationship between multi-modal predictions and model performance, and experimentally demonstrate the problems of existing distillation objectives in fitting multi-modal distributions. In response to the identified problems, we introduce the word-level ranking loss, which is based on Spearman’s rank correlation coefficient (SRCC), to optimize the degree of consistency in the order of peak predictions between the teacher and student models. In particular, we convert the learning of multi-modal distributions during the KD process into the learning of the top-k sampling Holtzman et al. (2020) order. Through ranking loss, we ensure excellent compatibility with existing distillation objectives while fully leveraging the fine-grained information between different categories predicted by two distribution peaks. Additionally, we verify and demonstrate through real-scenario experiments the impact of introducing ranking loss into KD on the learning ability of student models for multi-modal distributions.
Experimental results indicate that the quality of multi-modal predictions is closely related to the performance of the model, while existing distillation objectives lack the ability to fit multi-modal distributions effectively. Subsequently, our proposed method effectively enhances the student model’s learning ability to predict multi-modal distributions during the distillation process and exhibits good compatibility with existing distillation objectives. We also validate the ranking loss on diverse datasets from multiple tasks, showing significant improvements in downstream KD tasks.
@@ 2.1 KD of LLMs
Nowadays, many LLMs are no longer open source due to commercial and other considerations. Therefore, based on the open-source nature of the model, KD of LLMs is frequently categorized into white-box KD Gu et al. (2023); Tan et al. (2023); Wen et al. (2023); Ko et al. (2024); Wu et al. (2024a) for open-source LLMs Touvron et al. (2023); Yang et al. (2023); Jiang et al. (2024) and black-box KD Zhou et al. (2023); Chen et al. (2024) for closed-source LLMs Brown et al. (2020); OpenAI (2023).
@@ 2.2 White-Box KD Objectives
Unlike the above, Wu et al. (2024b) argue that the predicted distribution of LLMs does not meet the conditions of continuity and standard Gaussian distribution. They theoretically and practically demonstrate that KL and RKL actually share the same optimization objective in LLMs’ KD. Additionally, they point out that KL and RKL follow different optimization paths, with one fitting from the head part first and the other from the tail part first, and propose dynamically combining KL and RKL into Adaptive Kullback-Leiber (AKL).
Inspired by existing research, we aim at exploring a KD objective that yields better multi-modal learning ability. However, we first need to validate the following two questions:
@@ 3.1 Metrics
The purpose of enhancing the ability to learn multi-modal distributions is to make the peak predictions of student and teacher models closer. Therefore, we first need to define the criteria for evaluating the similarity of the peak predictions of two models. Since top-k sampling is the most common sampling strategy in language models, and peak predictions well correspond to the results of top-k sampling. Hence we convert the measurement of the consistency of multi-modal predictions into the measurement of the consistency of top-k sampling results.
@@ 3.3 Validation for Existing KD Objectives
Analysis Based on the results in Figure 3, existing distillation objectives show no significant disparity in their impact on the ability to learn multi-modal distributions. This further confirms Wu et al.’s (2024b) view that KL and RKL share the same optimization objective in KD of LLMs. But more importantly, even after 20 epochs of training, student models still exhibit deficiencies in learning multi-modal distributions under the existing distillation objectives. Therefore, further exploration is necessary to identify distillation metrics that can enhance the model’s capability to learn multi-modal distributions.
@@ 4 Method
In this work, we enhance the learning of peak predictions in KD of LLMs by introducing a new optimization objective of word-level ranking loss. The new optimization objective focuses on the prediction order of high probabilities between student and teacher models, enabling the student model to match the teacher model on critical predicted categories.
@@
Our approach allows that during the optimization process, the calculation of peak predictions is not limited to comparisons within a single category. As the red lines shown in Figure 4, $Q(i)$ needs to be compared with $Q(j)$ and $Q(k)$ based on the ranking position of $P(i)$ in the teacher’s predictions to minimize the ranking loss.
@@ 5.1 Baselines
Adaptive Kullback-Leiber divergence (AKL, Wu et al. 2024b) For the calculation process of AKL, we use the same experimental setup as in the original paper. We set the hyperparameter $\mu$ as 0.5 and the gap function $\epsilon(p(z),q(z))=|p(z)-q(z)|$.
@@ 5.2 Datasets and Models
databricks-dolly-15k Conover et al. (2023) A directive fine-tuning dataset covering various tasks. We randomly select 14,000 samples for the training set and 800 samples for the test set. We use ROUGE scores Lin (2004) as the evaluation metric to test the generative performance.
@@
For all KD tasks, we employ Llama-2-7B Touvron et al. (2023) as the teacher model and Tinyllama-1.1B Zhang et al. (2024) as the student model. Prior to distillation, we have fine-tuned the teacher models on the respective datasets to adapt it to the tasks. Except in the GSM8K task, we directly use gsm8k-rft-llama7b2-u13b model released by Yuan et al. (2023) due to its excellent performance.
@@ 5.3 Results in the Pre-Training Task
Table 2: Experimental results on test set of downstream tasks. "+R" represents that we have added an additional fixed-ratio ranking loss. "R-1", "R-2", and "R-L" are abbreviations for ROUGE-1, ROUGE-2, and ROUGE-L, respectively. We have also marked the improvement after combining the original objectives with ranking loss in parentheses for the most important metrics in the table.
Training With the same experimental setup as in Section 3.3, we test the effectiveness of these five distillation objectives combined with ranking loss. For ranking loss, we align predictions of the top 15 between student and teacher models. We employ a fixed ratio for loss allocation, as $\mathcal{L}_{\text{total}}=2\cdot\mathcal{L}_{\text{Ranking}}+\mathcal{L}_{\text{logits}}$. The choice of ranking ranges and the allocation of losses is discussed in Appendix D. For the convenience of subsequent ablation analysis, in addition to the multi-modal distribution similarity, we also evaluate the perplexity and CR of the top-1 prediction. The experimental results are presented in Table 1, while more detailed training information can be found in Appendix B. We also present extra experiments in Appendix E to verify the generalization ability of our method.
Results Table 1 shows the various metrics measured on the learned data for the student model before training and after KD training with different objectives. The results indicate that when combined with ranking loss, all five different objectives significantly improve the student model in terms of the multi-modal consistency metric with the teacher model during the distillation process. Moreover, while the similarity of the multi-modal distribution improves, the top-1 accuracy and perplexity performance are not negatively affected and even show slight improvements. To make the metrics for the similarity of multi-modal distributions more intuitive, we present visualized results in Figure 5, which show the improvement of ranking loss in a clearer manner. Compared to the mean scores of original objectives, our method improves the CR metric by approximately 30% to 95% across different ranges, and the OR metric is improved by about 50% to 120% across different ranges.
Overall, our approach significantly improves the efficiency of aligning multi-modal predictions between the student and teacher models during the distillation process. Furthermore, for the five different distillation objectives, ranking loss demonstrates stable and effective improvements, showcasing its excellent compatibility with all these commonly used KD objectives.
5.4 Results in Downstream Tasks
Although experiments in pre-training task fully validate that our method achieves the motivation of improving multi-modal prediction distribution learning during the KD process, its effectiveness in improving the performance of the student model after distillation still requires further verification. Therefore, we conduct thorough validation of our method on datasets from multiple different downstream tasks to demonstrate its contribution to improving the performance of the student model.
@@ 5.4 Results in Downstream Tasks
Results The experimental results show the scores of the student models after training with the baseline method and our method, where the highest scores for each metric of each task are achieved by our method. The table also shows the improvements after combining our proposed ranking loss with existing word-level distillation objectives. This combination enhances the performance of existing methods on nearly all metrics, with significant improvements on most. Specifically, after introducing the ranking loss, most of the accuracy improvements of the student model on the GSM8K test set are over 20% compared to original objective, most ROUGE-L scores on the Dolly test set improves by over 1.0 point, and most ROUGE-L score on the Xsum test set improves by over 0.7 points. Especially, when only using the ranking loss, our method learns only the peak predictions, which account for only about 0.0001% of the total categories, yet surpasses most existing distillation objectives in evaluations across multiple tasks. This not only demonstrates the importance of learning peak predictions, but also showcases the outstanding performance of our method in peak predictions alignment.
In summary, experiments on downstream tasks validate that our method significantly improves the performance of student model in the KD process. Compared to other optimization objectives of soft labels, our method demonstrates excellent competitiveness and compatibility. In addition, the consistent performance improvements across different datasets confirm the generality and robustness of our approach.
@@ 6.1 Ablation Study
In this section, we use further ablation analysis to reveal whether the performance improvement brought by our method is due to the enhanced ability of multi-modal distribution learning during the KD process.
Based on the respective results and analysis in Section 5.3 and 5.4, we can conclude a preliminary ablation conclusion that the improvement in the ability to learn multi-modal distributions and the enhancement in downstream task performance are indeed attributed to the introduction of ranking loss.
Furthermore, we can observe in Table 1 that although the improvement of the top prediction accuracy is modest by introducing the ranking loss, the peak prediction alignment at other positions is significantly improved. Therefore, ranking loss does not have a significant impact on the actual learning efficiency of the top-1 prediction.
However, in downstream tasks, we use a greedy decoding strategy, which should not exhibit better performance when there is no significant improvement in the consistency of the top prediction. In fact, the reason is that in downstream tasks, we make the student model learn task-related prediction patterns more efficiently by enhancing its ability to learn from multi-modal distributions. Therefore, our method achieve better performance with the same number of training steps.
Based on the above analysis, we can more confidently get the conclusion that ranking loss primarily improves the fitting of multi-modal distributions, with modest impact on the alignment of top-one predictions. Therefore, the existing objectives without the addition of ranking loss can be regarded as an ablation of the ability to learn multi-modal distributions, leading to worse results. This further shows the importance of aligning multi-modal predictive distributions in KD of LLMs.
@@ 6.2 Case Study
We conduct a case study within this interesting phenomenon in Appendix G to further analyze and demonstrate the importance of peak predictions learning.
@@ 7 Conclusion
In this paper, we propose ranking loss based knowledge distillation, a new objective function that improves the efficiency of aligning peak predictions of student and teacher models during white-box KD. We verify the importance of aligning multi-modal distributions through experiments and highlight the inefficiency of existing KD objectives in learning multi-modal distributions. Most importantly, we propose a word-level ranking loss to the existing KD objectives for more efficient alignment of multi-modal distributions. Our extensive experiments clearly demonstrate that our method effectively improves the multi-modal distribution alignment between teacher and student models, leading to significant performance gains in different downstream tasks.
@@
Acknowledgments
This work is supported by National Key R&D Program of China 2022ZD0160602 and the Natural Science Foundation of China 62122088.